In a pair of complementary angles one angle measures in (3x-40)degrees and the other is (4x+18)degrees what is the measure of the larger angle?
Found 2 solutions by Edwin McCravy, addingup:
Answer by Edwin McCravy(20077)

"Complementary angles" means that if you add their measures, 
you get 90°.

To add them you write them with a plus sign between them, like this:

     (3x-40)° + (4x+18)°

Then to say that when you add them, you get 90°, you put an equal 
sign after that, and then put 90° after that, like this:

     (3x-40)° + (4x+18)° = 90°

 Then you drop the ° marks and solve for x:

       (3x-40) + (4x+18) = 90
       3x - 40 + 4x + 18 = 90

                 7x - 22 = 90
                    + 22  +22
                ------------- 
                 7x     = 112
                      x = 112/7
                      x = 16

 The angle (3x-40)degrees is 
           (3·16-40)degrees or
           (48-40)degrees or
           8 degrees

 The angle (4x+18)degrees is 
           (4·16+18)degrees or
           (64+18)degrees or
           82 degrees

So the angles are 8° and 82°, which are complementary 
because when you add them you get 90°.

The larger one is 82°.

Edwin

Answer by addingup(3677)

3x-40+4x+18 = 90
7x-22 = 90
7x = 68
x = 9.71
------------------
One angle:
3(9.71)-40 = negative result, not possible
The other:
4(9.71)+18 = 56.84
If this second angle is correct, the first angle would be 90-56.81 = 33.19 because the property of complementary angles is that they add up to 90 degrees.